How many grams of water can be frozen at 0oC with the removal of -129.8 J? (Specific Heat = 4.2 𝐽/𝑔o𝐶, Enthalpy of fusion 6010 J/mol, Enthalpy of vaporization = 40700 J/mol)

1
Expert's answer
2021-04-19T05:14:42-0400

Number of moles frogen

=

6010 / 129.8


= 46.3 mol .


Mass of H2O = ( 46.3 × 18 ) g =

832 . 1538